3. A study by Consumer Reports showed that 64% of supermarket shoppers believe supermarket brands to be as good as national name brands. To investigate whether this result applies to its own store, the manager of the Publix in Clemson, SC asked a random sample of 100 shoppers whether they believed that supermarket brand ketchup was as good as the national brand ketchup, 52 believed the did believe the supermarket brand was as good as the national brand or not. Let a=10.
1
Expert's answer
2019-11-12T12:15:41-0500

Null hypothesis "H_0: p=0.64."

Alternative hypothesis "H_a: p\ne 0.64."

Test statistic: "z=\frac{0.52-0.64}{\sqrt{\frac{0.64(1-0.64)}{100}}}=-2.50."

P-value: "p=2P(Z<-2.50)=0.012."

Since the P-value is less than 0.1, reject the null hypothesis.

There is a sufficient evidence that the Consumer report result does not apply to

the Publix in Clemson, SC.